# Characterization of the chaotic phase in the tilted Bose-Hubbard model

### v1

@@ I Introduction
The study of chaotic and non-ergodic regimes in many-particle systems has received a lot of attention in recent years [8, 9, 10], given the importance of such phases for the possible engineering of a bespoke dynamical behaviour in an otherwise complex system. The identification of the chaotic phase is, e.g., crucial to ensure an adequate performance of current quantum computing architectures [11, 12, 13], and it also proves key for the benchmarking of quantum simulators [14, 15].
@@
Despite the significance of the tilted Bose-Hubbard Hamiltonian, its chaotic regime [66, 67, 68] remains unidentified. This work comes to fill this gap, unveiling the location and extension of the chaotic phase as functions of energy and model parameters, that will prove valuable for future theoretical and experimental studies of this system. Interestingly, our analysis reveals that that chaotic phase of the bare Bose-Hubbard model can in fact be enhanced by a moderate tilt.
@@ IV Energy-resolved chaotic phase
In order to ascertain the existence of a chaotic phase in the spectrum of the tilted Bose-Hubbard Hamiltonian [Eq. (1)], we calculate numerically the distribution of the level spacing ratios $r_{n}$ [Eq. (3)] and the generalized fractal dimension $\tilde{D}_{1}$ [Eq. (6)] from the whole spectrum of a system with $N=10$ bosons on $L=10$ sites as functions of the parameters $J$, $U$ and $F$. In this work, we restrict ourselves to lattices at unit filling ($N=L$), and the GFDs are calculated in the onsite Fock basis [Eq. (2)].
@@ V Emergence of chaos from the perspective of a homogeneous density configuration
By following the $E=0$ trajectory in the $(\varepsilon,J/U)$ and $(\varepsilon,F/J)$ spaces, one can see how the chaotic phase unfolds from the perspective of the state |$\ket{\boldsymbol{1}}$, hence identifying the range of parameter values where ergodicity could be experimentally observed dynamically with such initial state. The chaotic phase for the homogeneous Fock state as a function of $U/J$ and $F/J$ is shown in Fig. 6 for a system with $L=N=11$, where the usual figures of merit are numerically obtained by considering the $\approx 200$ eigenstates with energies closest to $E_{\ket{\boldsymbol{1}}}$. The ergodic domain is best recognizable from the behaviour of $\operatorname{var}\big{(}\tilde{D}_{1}\big{)}$, and roughly extends within the region $0.04\lesssim U/J\lesssim 4$ and $F/J\lesssim 1.5$, albeit the upper limit in $F$ increases with the interaction strength.

### v2

@@ I Introduction
The study of chaotic and non-ergodic regimes in many-particle systems has received a lot of attention in recent years [8, 9, 10, 11], given the importance of such phases for the possible engineering of a bespoke dynamical behaviour in an otherwise complex system. The identification of the chaotic phase is, e.g., crucial to ensure an adequate performance of current quantum computing architectures [12, 13, 14], and it also proves key for the benchmarking of quantum simulators [15, 16].
@@
Despite the significance of the tilted Bose-Hubbard Hamiltonian, its chaotic regime [67, 68, 69], which may be associated with unstable Bloch oscillation dynamics [70, 71], has not yet been fully identified. This work comes to fill this gap, unveiling the location and extension of the chaotic phase as functions of energy and model parameters, that will prove valuable for future theoretical and experimental studies of this system. Interestingly, our analysis reveals that that chaotic phase of the bare Bose-Hubbard model can in fact be enhanced by a moderate tilt.
@@ III Identification of quantum chaos
III.1 Numerical approach
To determine the emergence of quantum chaos, we calculate numerically the distribution of the level spacing ratios $r_{n}$ and the generalized fractal dimensions $\tilde{D}_{q}$ using exact diagonalization, making use in particular of the PETSc [102, 103, 104] and SLEPc [105] libraries, which permit an efficient parallelization of the numerical task to be run on High-Performance Computing facilities. A proper strategy using a combination of MPI and OpenMP [106] optimize the computational resources needed to obtain a few hundred eigensates in the excitation spectrum of our model for systems with a Hilbert space dimension of ${\mathcal{D}}\lesssim 5\times 10^{6}$. Additionally, the full spectrum of systems with ${\mathcal{D}}\lesssim 2\times 10^{5}$ can be tackled using the ScaLAPACK library [107].
@@ IV Energy-resolved chaotic phase
In order to ascertain the existence of a chaotic phase in the spectrum of the tilted Bose-Hubbard Hamiltonian [Eq. (1)], we first calculate the distribution of the level spacing ratios $r_{n}$ [Eq. (3)] and the generalized fractal dimension $\tilde{D}_{1}$ [Eq. (6)] from the whole spectrum of a system with $N=10$ bosons on $L=10$ sites as functions of the parameters $J$, $U$ and $F$. In this work, we restrict ourselves to lattices at unit filling ($N=L$), and the GFDs are calculated in the onsite Fock basis [Eq. (2)].
@@ V Emergence of chaos from the perspective of a homogeneous density configuration
By following the $E=0$ trajectory in the $(\varepsilon,J/U)$ and $(\varepsilon,F/J)$ spaces, one can see how the chaotic phase unfolds from the perspective of the state |$\ket{\boldsymbol{1}}$, hence identifying the range of parameter values where ergodicity could be experimentally observed dynamically with such initial state. The chaotic phase for the homogeneous Fock state as a function of $U/J$ and $F/J$ is shown in Fig. 6 for a system with $L=N=11$, where the usual figures of merit are numerically obtained by considering the $\approx 200$ eigenstates with energies closest to $E_{\ket{\boldsymbol{1}}}$. The ergodic domain is best recognizable from the behaviour of $\operatorname{var}\big{(}\tilde{D}_{1}\big{)}$, and roughly extends within the region $0.04\lesssim U/J\lesssim 4$ and $F/J\lesssim 1.5$, albeit the upper limit in $F$ increases with the interaction strength. It is also worth emphasizing that the observed chaotic phase lies within the region of unstable Bloch oscillation dynamics, identified from a mean-field approach in Refs. [70, 71], and demarcated by a dashed blue line in Fig. 6.
@@ Acknowledgements.
Data Availability Statement
The data to reproduce all figures in this article are openly available at http://hdl.handle.net/10366/164618.